There is a deck of 72 cards. 5 cards will be drawn. The order of the draw IS important. What is the number of possible 5 card draws?
1
Expert's answer
2020-07-29T14:41:27-0400

Given :

Total number of cards = 72

Use permutation when order is important.


Formula : when r items selected from n items , nPr= "\frac{n!}{(n-r)!}"


Solution :

72P5= "\frac{72!}{(72-5)!}"


= "\frac{72!}{67!}"


= 1678985280

Number of possible draws= 1678985280 ways